# CellCLIP – Learning Perturbation Effects in Cell Painting via Text-Guided Contrastive Learning

### v2

@@ 5.2 Intra-modality retrieval
We report results for these tasks in Table 3. We found that cross-modal CL outperformed our unimodal baselines. Moreover, within the considered CL approaches, we found that CellCLIP strongly outperformed all other approaches for replicate detection. For perturbation matching, when comparing perturbations targeting the same genes within the same perturbation class (e.g., compounds vs compounds), CellCLIP again performs the best among all approaches. For cross-class perturbation matching, (e.g. CRISPR vs compounds), CellCLIP again performs the best, though the overall performance for all machine-learning-based methods remains low. This suggests that, despite targeting the same gene, morphological changes across different perturbation classes remain highly distinct, aligning with previous findings [10].
@@
Table 3 reports recall of known biological relationships among genetic perturbations in RxRx3-core. Across all benchmark databases, we found that CellCLIP achieved the best recovery of known gene-gene relationships compared to baseline models. Notably, we found that replacing the CWCL loss used in CellCLIP with the standard CLIP loss leads to worse performance on this task (Table S.4), illustrating the benefits of using soft labeling for alignment in the image profile space $\mathcal{P}$. Altogether, these results further demonstrate that CellCLIP can recover meaningful relationships between perturbations in its image profile latent space.
@@ A.1 Effects of Different Image Profile Encoding
CellCLIP provides a flexible framework for integrating off-the-shelf pretrained image foundation models into Cell Painting analyses (Section 3.2). To understand the impact of different vision encoding backbones on CellCLIP’s performance, we conducted an ablation study where we varied CellCLIP’s image encoder while holding all other aspects of our framework constant. Specifically, for this experiment we considered DINOv1 along with DINOv2 models of varying sizes. Aligning with results for natural images, we found that increases in model size broadly led to increased performance on our retrieval tasks (Table S.1).
@@
A.2 Effect of Various Pooling Strategy
Table S.2 summarizes the retrieval performance across different pooling strategies as described in Section 3.1. Overall, attention-based pooling yields the best performance, particularly on the Recall@10. Among non-attention methods, mean pooling consistently outperforms median, K-means, and hierarchical clustering, suggesting its effectiveness as a simple yet strong baseline.
@@ A.2 Effect of Various Pooling Strategy
A.3 Effect of Different Contrastive Loss
Table S.3 summarizes cross-modal retrieval performance across different contrastive learning losses. CWCL consistently achieves the highest Recall@5 and Recall@10 in both perturbation-to-profile and profile-to-perturbation tasks, while remaining competitive at Recall@1. CLIP ranks second overall, and CLOOB shows the weakest performance. Although both CWCL and S2L use similarity-based weighting, CWCL outperforms S2L across all retrieval metrics. This suggests that computing similarity from embeddings extracted from pretrained imaging foundation models is more effective for alignment than using projected embeddings in the shared space, as defined in Fradkin et al. [16].
Table S.4 presents intra-modal gene–gene recovery performance. Sigmoid loss (SigCLIP) achieves the highest recall across all benchmarks, followed by CWCL. This indicates that contrastive objectives with relaxed pairwise penalties may better preserve biological structure. Overall, CWCL demonstrates strong performance across both settings, highlighting its ability to balance cross-modal alignment with intra-modal consistency.
@@ A.3 Effect of Different Contrastive Loss
A.4 Zero-shot Biological Relationship Recovery - RxRx3-core
We evaluate recall over the most extreme 2% to 20% of the similarity distribution on RxRx3-core (Figure S.1), including pathway annotations from CORUM, HuMAP, Reactome, SIGNOR, and STRING databases. As shown in the figure below, CellCLIP consistently achieves the highest recall across thresholds, outperforming existing multimodal contrastive methods such as CLOOME and MolPhenix. Among CellCLIP variants, performance is comparable across different image tokenizers, indicating robustness to the choice of visual backbone (Table S.5).
@@ Appendix B CellCLIP Training & Implementation Details
To generate channel profiles $z_{k}$ (Section 3.1), we experimented with a range of image foundation models, including DINO, DINOv2 (small, base, large, giant), and CA-MAE (OpenPhenom-S/16) (Table S.6). All model checkpoints were obtained from Hugging Face. Cell Painting images were treated as grayscale inputs, preprocessed according to the requirements of each model, and augmented using a multi-crop strategy to enhance robustness.
@@ C.1 Cross-Modal Retrieval
[…]
@@ C.2.3 Zero-shot Recovery of Known Biological Relationship (RxRx3-core [27])
This task evaluates whether the learned embeddings capture biologically meaningful structure by assessing their ability to recover known functional relationships among perturbations, such as genes within the same pathway or protein complex. Pathway annotations are obtained from HuMAP, CORUM, StringDB, Reactome, and SIGNOR (Table S.7).
For each perturbation $i$, we compute an aggregated embedding $\tilde{p}_{i}$ as described in Section C.2.2 but using the median for $\mathcal{A}$ as in Kraus et al. [26]. For each perturbation $i$ we then evaluate the cosine similarity of $\tilde{p}_{i}$ with embeddings $\tilde{p}_{j}$ from other perturbations. All perturbation pairs are then ranked by similarity to assess the recovery of known biological relationships.
@@
A random embedding space yields a baseline recall of 10%. Higher recall indicates stronger alignment with curated biological interaction networks. We also report recall@2% ([0.01, 0.99]) to recall@ 20% ([0.10, 0.90]) in Figure S.1. Evaluation was performed following the official RxRx3-core benchmark protocol. [^4] EFAAR Benchmark
@@ Appendix D Datasets & Preprocessing
Sanchez-Fernandez et al. [37] refined the dataset by removing images that were out of focus, exhibited high fluorescence, or contained untreated control cells. The final dataset comprises 759,782 images linked to 30,404 unique molecules, split into training (674,357), validation (28,632), and test (56,793) sets. The processed dataset is publicly available at [^7] Curated dataset of Bray et al. [4]. Next, for retrieval evaluation on unseen compounds Section 4.1, following Sanchez-Fernandez et al. [37], we removed samples from the test set that corresponded to the same molecule and plate to mitigate plate effects. The remaining samples are referred to as the final "test set", which consists of 2,115 compounds.
@@
CP-JUMP1 [10] comprises 186,925 nine-channel (3 bright field channels) microscopy images of resolution 1080 $\times$ 1080. This dataset features a comprehensive collection of perturbations conducted on U2OS and A549 cell lines, including 52 replicates. For perturbations, it includes 301 small-molecule compounds ( 46 controls), 335 sgRNAs (CRISPR) targeting 175 genes ( 88 control sgRNAs), and 175 ORFs (45 controls) for the corresponding genes. [10] also provides annotations of associations between genes and compounds, enhancing its utility for exploring gene function and compound effects. We split the dataset into 70/10/20 for training, validation, and testing. Raw data, relevant metadata, and gene annotation can be found in [^8] Github repository of Chandrasekaran et al. [10]
@@
RxRx3-core[27] [^9] RxRx3-core dataset is a curated subset of the RxRx3 dataset [15]. It includes 222,601 six-channel fluorescent microscopy images of HUVEC cells stained using a modified Cell Painting protocol, covering 736 gene knockouts (non-blinded) and 1,674 compounds across eight concentrations, along with control wells. Gene-gene relationships are benchmarked using Reactome, HuMAP, SIGNOR, StringDB, and the CORUM collection, Table S.7. Additional details regarding RxRx3-core can be found in Kraus et al. [27].
@@ E.2 Hyperparameter Search
For CellCLIP, CLOOME (Phenomic), and MolPhenix*, we perform hyperparameter search over the following ranges: learning rate $\{5\times 10^{-5},1\times 10^{-5},1\times 10^{-4},2\times 10^{-4},3\times 10^{-4},5\times 10^{-4},1\times 10^{-3},1\times 10^{-2}\}$, warmup steps $\{200,500,800,1000,1500\}$, number of learning rate restarts $\{1,2,3,4\}$, and temperature $\tau\in\{14.3,20,30\}$. For batch size, we use the largest possible setting on 8 RTX 6000 GPUs: CellCLIP, MolPhenix*, and CLOOME (Phenomic) are trained with a batch size of 512, 768, and 768, while CLOOME is trained with a batch size of 256.

### v3

@@ 5.2 Intra-modality retrieval
We report results for these tasks in Table˜4. We found that cross-modal CL outperformed our unimodal baselines. Moreover, within the considered CL approaches, we found that CellCLIP strongly outperformed all other approaches for replicate detection. For perturbation matching, when comparing perturbations targeting the same genes within the same perturbation class (e.g., compounds vs compounds), CellCLIP again performs the best among all approaches. For cross-class perturbation matching, (e.g. CRISPR vs compounds), CellCLIP again performs the best, though the overall performance for all machine-learning-based methods remains low. This suggests that, despite targeting the same gene, morphological changes across different perturbation classes remain highly distinct, aligning with previous findings [10].
@@
Table˜4 reports recall of known biological relationships among genetic perturbations in RxRx3-core. Across all benchmark databases, we found that CellCLIP achieved the best recovery of known gene-gene relationships compared to baseline models. Notably, we found that replacing the CWCL loss used in CellCLIP with the standard CLIP loss leads to worse performance on this task (Table˜S.5), illustrating the benefits of using soft labeling for alignment in the image profile space $\mathcal{P}$. Altogether, these results further demonstrate that CellCLIP can recover meaningful relationships between perturbations in its image profile latent space.
@@ A.1 Effects of Different Image Profile Encoding
CellCLIP provides a flexible framework for integrating off-the-shelf pretrained image foundation models into Cell Painting analyses (Section˜3.2). To understand the impact of different vision encoding backbones on CellCLIP’s performance, we conducted an ablation study where we varied CellCLIP’s image encoder while holding all other aspects of our framework constant. Specifically, for this experiment we considered DINOv1 along with DINOv2 models of varying sizes. Aligning with results for natural images, we found that increases in model size broadly led to increased performance on our retrieval tasks (Table˜S.1).
@@
A.2 Effect of Various Pooling Strategy
Table˜S.2 summarizes the retrieval performance across different pooling strategies as described in Section˜3.1. Overall, attention-based pooling yields the best performance, particularly on the Recall@10. Among non-attention methods, mean pooling consistently outperforms median, K-means, and hierarchical clustering, suggesting its effectiveness as a simple yet strong baseline.
A.3 Impact of top-$n$ instances
Table S.3: Retrieval performance after removing the top-$n$ contributing instances. Results are reported as Recall@$k$ (%) with $k=1,5,10$.
Removal of top-$n$ instances Perturb-to-profile (%) Profile-to-perturb (%)
R@1 R@5 R@10 R@1 R@5 R@10
0 1.32 4.59 7.51 1.18 4.82 7.56
1 1.08 4.60 7.13 1.04 4.58 7.61
2 1.13 4.25 6.38 0.99 4.34 6.47
3 1.13 4.20 5.91 0.89 4.06 6.09
4 0.04 2.50 5.01 0.09 3.76 5.34
5 0.00 1.46 3.68 0.00 2.88 4.34
To mitigate the influence of false positives, we applied per-perturbation pooling of cellular profiles using the ABMIL pooling operator [24], which adaptively attends to samples most representative of a perturbation. We further investigated the importance of accounting for false positives by removing the most highly attended samples (i.e., those most indicative of effective perturbation). cross-modal retrieval tasks. This ablation led to substantial performance degradation (see Table˜S.3), suggesting that for each perturbation, only a relatively small fraction of cells exhibit strong perturbation effects.
A.4 Effect of Different Contrastive Loss
Table˜S.4 summarizes cross-modal retrieval performance across different contrastive learning losses. CWCL consistently achieves the highest Recall@5 and Recall@10 in both perturbation-to-profile and profile-to-perturbation tasks, while remaining competitive at Recall@1. CLIP ranks second overall, and CLOOB shows the weakest performance. Although both CWCL and S2L use similarity-based weighting, CWCL outperforms S2L across all retrieval metrics. This suggests that computing similarity from embeddings extracted from pretrained imaging foundation models is more effective for alignment than using projected embeddings in the shared space, as defined in Fradkin et al. [16].
Table˜S.5 presents intra-modal gene–gene recovery performance. Sigmoid loss (SigCLIP) achieves the highest recall across all benchmarks, followed by CWCL. This indicates that contrastive objectives with relaxed pairwise penalties may better preserve biological structure. Overall, CWCL demonstrates strong performance across both settings, highlighting its ability to balance cross-modal alignment with intra-modal consistency.
@@ A.4 Effect of Different Contrastive Loss
A.5 Zero-shot Biological Relationship Recovery - RxRx3-core
We evaluate recall over the most extreme 2% to 20% of the similarity distribution on RxRx3-core (Figure S.1), including pathway annotations from CORUM, HuMAP, Reactome, SIGNOR, and STRING databases. As shown in the figure below, CellCLIP consistently achieves the highest recall across thresholds, outperforming existing multimodal contrastive methods such as CLOOME and MolPhenix. Among CellCLIP variants, performance is comparable across different image tokenizers, indicating robustness to the choice of visual backbone (Table˜S.6).
@@ A.5 Zero-shot Biological Relationship Recovery - RxRx3-core
Table S.7: Retrieval performance with different BERT parameter sizes. Results are reported as Recall@$k$ (%) with $k=1,5,10$.
# BERT params Perturb-to-profile (%) Profile-to-perturb (%)
R@1 R@5 R@10 R@1 R@5 R@10
110M (ours) $1.18\pm 0.20$ $4.49\pm 0.06$ $7.37\pm 0.20$ $1.25\pm 0.10$ $4.82\pm 0.10$ $7.39\pm 0.23$
64M $1.53\pm 0.17$ $4.40\pm 0.21$ $6.90\pm 0.18$ $1.41\pm 0.17$ $4.67\pm 0.24$ $6.98\pm 0.14$
11.3M $1.13\pm 0.14$ $4.04\pm 0.31$ $6.28\pm 0.19$ $1.28\pm 0.17$ $4.11\pm 0.24$ $6.20\pm 0.19$
A.6 Effect of Text Encoder Parameters
In Table˜S.7, we present additional experiments using smaller BERT variants (64M and 11.3M parameters) from Turc et al. [44] as the text encoder for CellCLIP on the cross-modal retrieval task [4]. For reference, the MPNN+ graph neural network encoder in our MolPhenix baseline contains approximately 10M parameters. While performance decreases slightly with smaller BERT models, even the 11.3M-parameter variant substantially outperforms the MPNN+ baseline. These results indicate that the gains in cross-modal retrieval are not merely due to larger model capacity, but rather arise from the effectiveness of our text-based encoding strategy.
Table S.8: Retrieval performance with different perturbation encoders. Results are reported as Recall@$k$ (%) with $k=1,5,10$.
Perturb. Encoder Perturb-to-profile (%) Profile-to-perturb (%)
R@1 R@5 R@10 R@1 R@5 R@10
BiomedBERT $1.46\pm 0.10$ $4.11\pm 0.13$ $6.19\pm 0.26$ $1.51\pm 0.09$ $4.34\pm 0.13$ $6.57\pm 0.19$
KV-PLM $1.29\pm 0.19$ $4.41\pm 0.31$ $6.92\pm 0.32$ $1.41\pm 0.16$ $4.57\pm 0.30$ $6.81\pm 0.21$
BERT $1.18\pm 0.20$ $4.49\pm 0.06$ $7.37\pm 0.20$ $1.25\pm 0.10$ $4.82\pm 0.10$ $7.39\pm 0.23$
A.7 Do Domain-Adapted BERT Help?
In Table˜S.8, we compare several pretrained BERT models, including BiomedBERT [49], trained on biomedical PubMed corpora, and KV-PLM [47], a BERT-based model pretrained on both biomedical texts and SMILES strings. Interestingly, while KV-PLM outperforms BiomedBERT, it still underperforms relative to our original BERT choice, suggesting that domain adaptation alone does not guarantee superior performance for cross-modal retrieval.
@@ Appendix B CellCLIP Training & Implementation Details
To generate channel profiles $z_{k}$ (Section˜3.1), we experimented with a range of image foundation models, including DINO, DINOv2 (small, base, large, giant), and CA-MAE (OpenPhenom-S/16) (Table˜S.9). All model checkpoints were obtained from Hugging Face. Cell Painting images were treated as grayscale inputs, preprocessed according to the requirements of each model, and augmented using a multi-crop strategy to enhance robustness.
@@ C.1 Cross-Modal Retrieval
$\text{Recall@k}_{\mathcal{U}\rightarrow\mathcal{V}}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}\left[\text{rank}_{\mathcal{V}}(q_{i}\mid p_{i})\leq k\right],$ (11)
@@ C.2.3 Zero-shot Recovery of Known Biological Relationship (RxRx3-core [28])
This task evaluates whether the learned embeddings capture biologically meaningful structure by assessing their ability to recover known functional relationships among perturbations, such as genes within the same pathway or protein complex. Pathway annotations are obtained from HuMAP, CORUM, StringDB, Reactome, and SIGNOR (Table˜S.10).
For each perturbation $i$, we compute an aggregated embedding $\tilde{p}_{i}$ as described in Section˜C.2.2 but using the median for $\mathcal{A}$ as in Kraus et al. [27]. For each perturbation $i$ we then evaluate the cosine similarity of $\tilde{p}_{i}$ with embeddings $\tilde{p}_{j}$ from other perturbations. All perturbation pairs are then ranked by similarity to assess the recovery of known biological relationships.
@@
A random embedding space yields a baseline recall of 10%. Higher recall indicates stronger alignment with curated biological interaction networks. We also report recall@2% ([0.01, 0.99]) to recall@ 20% ([0.10, 0.90]) in Figure˜S.1. Evaluation was performed following the official RxRx3-core benchmark protocol. [^4] EFAAR Benchmark
@@ Appendix D Datasets & Preprocessing
Sanchez-Fernandez et al. [38] refined the dataset by removing images that were out of focus, exhibited high fluorescence, or contained untreated control cells. The final dataset comprises 28.4k images linked to 10.8k unique molecules, split into training, validation, and test sets. The processed dataset is publicly available at [^7] Curated dataset of Bray et al. [4]. Next, for retrieval evaluation on unseen compounds Section˜4.1, following Sanchez-Fernandez et al. [38], we removed samples from the test set that corresponded to the same molecule and plate to mitigate plate effects. The remaining samples are referred to as the final "test set", which consists of 2,115 unique compounds.
@@
CP-JUMP1 [10] comprises approximately 340k eight-channel (3 bright field channels) microscopy images of resolution 1080 $\times$ 1080. We subsampled images per perturbation, resulting in a dataset of 186,925 images. This dataset features a comprehensive collection of perturbations conducted on U2OS and A549 cell lines, including 52 replicates. For perturbations, it includes 301 small-molecule compounds ( 46 controls), 335 sgRNAs (CRISPR) targeting 175 genes ( 88 control sgRNAs), and 175 ORFs (45 controls) for the corresponding genes. [10] also provides annotations of associations between genes and compounds, enhancing its utility for exploring gene function and compound effects. We split the dataset into 70/10/20 for training, validation, and testing. Raw data, relevant metadata, and gene annotation can be found in [^8] Github repository of Chandrasekaran et al. [10]
@@
RxRx3-core[28] [^9] RxRx3-core dataset is a curated subset of the RxRx3 dataset [15]. It includes 222,601 six-channel fluorescent microscopy images of HUVEC cells stained using a modified Cell Painting protocol, covering 736 gene knockouts (non-blinded) and 1,674 compounds across eight concentrations, along with control wells. Gene-gene relationships are benchmarked using Reactome, HuMAP, SIGNOR, StringDB, and the CORUM collection, Table˜S.10. Additional details regarding RxRx3-core can be found in Kraus et al. [28].
@@ E.2 Hyperparameter Search
For CellCLIP, CLOOME (Phenomic), and MolPhenix* results in Table˜1, we perform hyperparameter search over the following ranges: learning rate $\{5\times 10^{-5},1\times 10^{-5},1\times 10^{-4},2\times 10^{-4},3\times 10^{-4},5\times 10^{-4},1\times 10^{-3},1\times 10^{-2}\}$, warmup steps $\{200,500,800,1000,1500\}$, number of learning rate restarts $\{1,2,3,4\}$, and temperature $\tau\in\{14.3,20,30\}$. For batch size, we use the largest possible setting on 8 RTX 6000 GPUs: CellCLIP, MolPhenix*, and CLOOME (Phenomic) are trained with a batch size of 512, 768, and 768, while CLOOME is trained with a batch size of 256.
For CellCLIP ablation results in Table˜2, we performed a hyperparameter sweep over: Learning rate: $\{0.0005,0.001,0.005\}$, cosine annealing restart cycles: $\{5,6,7,8,9,10\}$, temperature $\tau$: $\{14.3,20,30\}$, epochs: $\{50,60,70\}$.
[…]
For both of the unimodal CA‑MAE and OpenPhenom‑S/16 models specifically, as these models’ training data already included CP-JUMP1, we did not perform any additional training steps with these models